OpenStudy (anonymous):

describe how to draw a line that passes through the origin and has a slope of -2/3?

OpenStudy (alyssa_xo):

use point slope form: \(y-y_1=m(x-x_1)\) where \((x_1,y_1)\) is the point we want the line to pass through and m is the slope

OpenStudy (alyssa_xo):

in our case it'd be \(y-0=\frac{-2}{3}(x-0)\)